OpenStudy (anonymous):

Solve the equation for x: 2x+1=5/(x+2).

OpenStudy (anonymous):

x=-3 and x=1/2

OpenStudy (anonymous):

(2x+1)(x+2)=5 2x^2+4x+x+2=5 2x^2+5x-3=0 now you can use the quadratic formula to get the roots.

OpenStudy (anonymous):

Danke!

OpenStudy (anonymous):

Bitte!